One card is drawn at random from a well-shuffled pack of 52 playing cards and the colour is noted(red or black).The card is returned to the pack and a second card is drawn and the colour noted. What is the probability of drawing:
a. two red cards
b. a red card followed by a black card
c. one card each of red and black?

Answer by greenestamps(13330)



(a) both red: P(R,R) = (1/2)(1/2) = 1/4

(b) red, then black: P(R,B) = (1/2)(1/2) = 1/4

(c) one red and one black (red, then black; or black, then red): P(R,B)+P(B,R) = 1/4+1/4 = 1/2